# A new economic and financial theory of money

### v5

@@ XX Discussion and Conclusions
The very good news was given in Sec. XVII. This section discussed the interactions of economic sub-systems and sovereign states. It started by describing how the economic system or global community accommodates multi-objective and multi-party optimization. This led to the concept of the geometry and topology of the economic system that is captured by the metric of exchange rates, $\Xi_{ij}$ given the discrete set of basis vectors $\left|\xi_{i}\right>$ forming a low dimensional manifold on a Hilbert space. A natural evolution first will lead to a separation of topology into two disconnected network graph clusters, one of the aesthetic sub-economies and another of the unaesthetic sub-economies. Whether it by migration of the members of the sub-economy, if there is enough freedom, or by replacement of the leaders of the unaesthetic sub-economies, usually by armed revolt; the economic system will then evolve to one aesthetic cluster.
History has shown us that the unaesthetic has dominated, though. There is a paradox of tyranny. The one person that must have an aesthetic ethos is the leader. The paradox arises from the fact that the leader is the one person who will benefit from an unaesthetic ethos. Unless a super majority of the community is educated, the leader can propagandize enough of the community to follow the greedy path of exploitation, so that the leader can stay in power. The leader can also change the rules so that the leader will always win the game. So, education, that is enlightenment, is fundamental for evolution to an aesthetic society. The additional fundamental ingredients are electronic currencies and AI based monetary, investment, and operational control.
@@
Instead, society has capitulated to the forces of external exploitation (ugliness) by building that exploitation (resistivity) into the dynamics of the sub-system (via debt financing), destroying the topological beauty of nature. This is the cause of both the existential angst of Jean-Paul Sartre in Les Jeux Sont Faits, and the profound acquiescence to fate in Latin American literature of authors such as Gabriel Garcia Márquez in Cien Años de Soledad.
We leave you with a final thought. Our current economic system is one large resistor. It needs to be transformed into a much more “conservative system” (that is Hamiltonian system) with significant savings and storage (that is potential energy). Creation of this capacity for storing economic energy will take significant financing. It can not and will not be financed with resistive financing (that is debt). A new method of “conservative financing” must be created (that is electronic currency). That is what this paper is proposing – true economic conservatism.
@@ Acknowledgements.
First of all, we would like to thank Ted Frankel and Michael Freedman and for giving a deep appreciation for the primal importance of geometry and topology to the sciences, including economics. To the Santa Fe Institute and Murray Gell-Mann, for a stay where the richness of complex systems and how to approach them from a scientific perspective was learned. To the Institut des Hautes Etudes Scientifique and Stephane Mallat, for another stay where the methods of topological analysis were inspired. Last, but not least, to Robert Baird who gave a deep love and understanding of economics.

### v6

@@ XXI Discussion and Conclusions
The very good news was given in Sec. XVIII. This section discussed the interactions of economic sub-systems and sovereign states. It started by describing how the economic system or global community accommodates multi-objective and multi-party optimization. This led to the concept of the geometry and topology of the economic system that is captured by the metric of exchange rates, $\Xi_{ij}$ given the discrete set of basis vectors $\left|\xi_{i}\right>$ forming a low dimensional manifold on a Hilbert space. A natural evolution first will lead to a separation of topology into two disconnected network graph clusters, one of the aesthetic sub-economies and another of the unaesthetic sub-economies. Whether by migration of the members of the sub-economy, if there is enough freedom, or by replacement of the leaders of the unaesthetic sub-economies, usually by armed revolt; the economic system will then evolve to one aesthetic cluster.
History has shown us that the unaesthetic has dominated, though. There is a paradox of tyranny. The one person that must have an aesthetic ethos is the leader. The paradox arises from the fact that the leader is the one person who will benefit from an unaesthetic ethos. Unless a super majority of the community is educated, the leader can propagandize enough of the community to follow the greedy path of exploitation, so that the leader can stay in power. The leader can also change the rules so that the leader will always win the game. So, education, that is enlightenment, is fundamental for evolution to an aesthetic society. The additional fundamental ingredients are electronic currencies and genAI based monetary, investment, and operational control.
@@
Instead, society has capitulated to the forces of external exploitation (ugliness) by building that exploitation (resistivity) into the dynamics of the sub-system (via debt financing), destroying the topological beauty of nature. This is the cause of both the existential angst of Jean-Paul Sartre in “Les Jeux Sont Faits” (Sartre, 1947), and the profound acquiescence to fate in Latin American literature of authors such as Gabriel Garcia Márquez in “Cien Años de Soledad” (Márquez, 1967).
In summary, this paper presented a new unified economic theory that relaxed the capitalistic diffusive approximation made for control so that there is coordinated, yet distributed, monetary and fiscal (print/grant investment) policy made to maximize GDP, not DCF. This is a socialist, not a capitalistic metric, where the economic system is treated as a web (multiscale graph) of economic communes. The local microeconomics is treated like the global macroeconomics. The viscous forecasting and control is replaced by a Generative Artificial Intelligence (genAI) forecasting and control. The control is distributed (that is, democratic), not centralized (that is autocratic). Monetarism argues for only an autocratic (centralized) capitalistic (profit based) monetary control, and Keynesianism argues for only an autocratic capitalistic fiscal (borrow/spend investment) control. This new theory argues for coordinated democratic socialist monetary and fiscal control. It does have philosophical connections to African Ubuntu, the Economics of Mutuality, Modern Monetary Theory, Marxian Economics and religious philosophy. It also makes the connection between genAI and the physics of economic collective systems. The result is an economic system that maximizes the good of society, not an economic system that maximally exploits society.
We leave you with a final thought. Our current economic system is one large resistor. It needs to be transformed into a much more “conservative system” (that is Hamiltonian system), an economic superconductor, with significant savings and storage (that is potential energy). Creation of this capacity for storing economic energy will take significant financing. It can not and will not be financed with resistive financing (that is debt). A new method of “conservative financing” must be created (that is transactional equity based Ubuntu Financing). That is what this paper is proposing – true economic conservatism.
@@ Acknowledgements.
First of all, we would like to thank Ted Frankel and Michael Freedman and for giving a deep appreciation for the primal importance of geometry and topology to the sciences, including economics. To the Santa Fe Institute and Murray Gell-Mann, for a stay where the richness of complex systems and how to approach them from a scientific perspective was learned. To the Institut des Hautes Etudes Scientifique and Stephane Mallat, for another stay where the methods of topological analysis were inspired. To Rev. Benjamin Larzelere III for intriguing and useful discussions about religious philosophy. Last, but not least, to Robert Baird who gave a deep love and understanding of economics.